OpenStudy (anonymous):

A 5-meter ladder leans against a wall. Suppose that the top is sliding down the wall at a rate of 1.2 m/s. calculate dx/dt when h=3m.

OpenStudy (anonymous):

key equation pythagorean theorem \[a^2+b^2=c^2\]

OpenStudy (anonymous):

if that still doesnt help, i suggest drawing a picture

OpenStudy (anonymous):

so it would be \[a ^{2} +3^{2} =5^{2}\]?? and then take the derivatives?

OpenStudy (anonymous):

nope dont substitute the 3 in yet

OpenStudy (anonymous):



OpenStudy (anonymous):

x^2+y^2=5^2 move the y or x to the other side x^2=25-y^2 then you take the derivative with respect to time dx/dt would be the rate x is increasing dy/dt would be the rate y is increasing

OpenStudy (anonymous):

anyways if you take the derivative with respect to time for x^2=25-y^2 what would you get?

OpenStudy (anonymous):

2x=25-2y

OpenStudy (anonymous):

?

OpenStudy (anonymous):

close, but you're taking the derivative with respect to t x and y are lengths, you would get 2x(dx/dt)=-2y(dy/dy) also the derivative of a constant is 0, so there wouldnt be a 25

OpenStudy (anonymous):

what we know top is sliding down at a rate of 1.2 m/s so dy/dt =-1.2 m/s we want to solve for dx/dt when h=3 so y=h=3 you can determine x from pythagorean theorem so all you need to do is solve for dx/dt